Gamer Co. has no debt. Its cost of capital is 9.8 percent. Suppose the company converts to a debt–equity ratio of 1. The interest rate on the debt is 6.9 percent. Ignore taxes for this problem.

What is the company’s new cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Cost of equity : ____%

What is its new WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

WACC: _____%  
Answer #1

1)

New cost of equity = Cost of unlevered equity + (Cost of unlevered equity - cost of debt)debt-equity ratio

New cost of equity = 9.8% + (9.8% - 6.9%)1

New cost of equity = 9.8% + 2.9%

New cost of equity = 12.70%

2)

Weight of equity = 1 / (1 + D/E)

Weight of equity = 1 / (1 + 1)

Weight of equity = 0.5

Weight of debt = 1 - 0.5 = 0.5

WACC = 0.5*0.127 + 0.5*0.069

WACC = 0.0635 + 0.0345

WACC = 0.0980 or 9.80%